OpenStudy (anonymous):

if y=3x+4 were changed to y=5x+4 how would the new function of the graph of the new function compare with the first one a.it would be less steep b. it would be steeper c. it would be shifted up d. it would be shifted right

OpenStudy (anonymous):

I recommend plotting these on the same XY plane. That should answer your question.

OpenStudy (anonymous):

So it would be a?

OpenStudy (abdullah1995):

hey did you find an answer yet ?

OpenStudy (anonymous):

no

OpenStudy (abdullah1995):

y=mx+b

OpenStudy (alexandervonhumboldt2):

look at slope @jessica1609 . y=ax+b where a is slope. you do see that second function has bigger slope. thus it is more steep,

OpenStudy (anonymous):

so b?

OpenStudy (abdullah1995):

@AlexandervonHumboldt2 explained it perfectly

OpenStudy (alexandervonhumboldt2):

yes B

OpenStudy (alexandervonhumboldt2):

thx @abdullah1995

OpenStudy (anonymous):

thank you @AlexandervonHumboldt2

OpenStudy (alexandervonhumboldt2):

np

OpenStudy (anonymous):

thank you guys

OpenStudy (alexandervonhumboldt2):

\(\huge\color{lime}{{\rm your}\rm~welcome!!!!}\) \(\Huge \color{blue}{\star^{ \star^{\star:)}}}\Huge \color{red}{\star^{ \star^{\star:)}}}\) \(\Huge \color{green}{\star^{ \star^{\star:)}}}\Huge \color{yellow}{\star^{ \star^{\star:)}}}\)

OpenStudy (anonymous):

what about when theres a question that says "the graph of a line gets __________ as the value of the slope gets bigger. a. less steep b.shorter c.steeper d. longer @AlexandervonHumboldt2

OpenStudy (alexandervonhumboldt2):

the bigger the slope, the steeper the line

OpenStudy (alexandervonhumboldt2):

the steeper the line, the bigger the slope

OpenStudy (anonymous):

so the line gets steeper thanks i understood @AlexandervonHumboldt2

OpenStudy (alexandervonhumboldt2):

yeah

OpenStudy (anonymous):

thank you @AlexandervonHumboldt2

OpenStudy (anonymous):

http://web.geogebra.org/app/ This website helps a lot when working with linear equations. Or graphing. Press Geometry.

OpenStudy (anonymous):

if you wanted to make the graph of y=2x+5 steeper which equation could you use? a. y=x+5 b. y=2x+8 c. y=4x+5 d. y=-2x+5 @AlexandervonHumboldt2

OpenStudy (alexandervonhumboldt2):

you need the biger slope. which euqtion have bigger slope than the original equation?

OpenStudy (anonymous):

b

OpenStudy (alexandervonhumboldt2):

no it has the same.

OpenStudy (anonymous):

so the slope is in the front

OpenStudy (anonymous):

like the front number

OpenStudy (alexandervonhumboldt2):

yes the front is lsope.

OpenStudy (anonymous):

thank you

OpenStudy (alexandervonhumboldt2):

np